# Stabilizing Thompson Sampling with Point Null Bayesian Response-Adaptive Randomization

### v1

@@ 3.1 Two group comparisons
Assume that the data are summarized by $y=\{\hat{\theta},\sigma\}$, where $\hat{\theta}$ is an effect estimate of the true treatment effect $\theta$ that quantifies the effect of a treatment over the control and $\sigma$ is the standard error of the estimate. For example, $\hat{\theta}$ could be an estimated (standardized) mean difference, log odds/rate/hazard ratio, risk difference, or regression coefficient. Suppose further that the estimate is (at least approximately) normally distributed, i.e., $\hat{\theta}\mid\theta\sim\mathrm{N}(\theta,\sigma^{2})$.
If the effect $\theta$ is oriented such that a positive effect indicates treatment benefit, the three hypotheses from Section 2.1 translate into:
$H_{-}\colon\theta<0$ versus $H_{0}\colon\theta=0$ versus $H_{+}\colon\theta>0$
@@
The point null hypothesis $H_{0}$ is a simple hypothesis with no free parameters, or equivalently, a point (Dirac) prior at zero. The $H_{-}$ and $H_{+}$ hypotheses are composite hypotheses and require prior distributions for the effect $\theta$. A natural choice is a $\theta\sim\mathrm{N}(\mu,\tau^{2})$ normal distribution whose support is truncated to the negative and positive side, respectively. This distribution may be specified based on prior knowledge (e.g., previous studies). In the absence of prior knowledge, it seems sensible to center the prior on zero ($\mu=0$) to represent clinical equipoise (Freedman,, 1987), as a zero-centered prior gives equal probability to harmful and beneficial effects. Averaging the prior over the three hypotheses leads to a spike-and-slab prior, as illustrated in Figure 2.
@@
To compute posterior hypothesis probabilities, specification of prior hypothesis probabilities is required. The prior probability of the null hypothesis $\Pr(H_{0})$ represents the a priori plausibility of an absent effect, and also acts as a tuning parameter that controls the degree of variability of RAR. An intuitive default is $\Pr(H_{0})=0.5$ as it represents the equipoise position of equal probability of an absent effect relative to a present (either harmful or beneficial) effect (Johnson,, 2013). For a given $\Pr(H_{0})$, it is then natural to set the prior probabilities of the other two hypotheses to $\Pr(H_{+})=\{1-\Pr(H_{0})\}\times\Phi(\mu/\tau)$ and $\Pr(H_{-})=\{1-\Pr(H_{0})\}\times\Phi(-\mu/\tau)$ so that to the prior distribution averaged over $H_{-}$ and $H_{+}$ is again the $\mathrm{N}(\mu,\tau^{2})$ normal distribution that was truncated in the first place. For example, when setting $\Pr(H_{0})=0.5$ and specifying a zero-centered prior ($\mu=0$) as in Figure 2, we obtain $\Pr(H_{+})=\Pr(H_{-})=0.5\times 0.5=0.25.$
With the prior densities and prior hypothesis probabilities specified, we can compute the marginal likelihood of the observed effect estimate under each hypothesis, and in turn obtain posterior probabilities (1). In the conjugate normal likelihood and prior framework, all of them can be straightforwardly derived in closed-form. Denoting by $\mathrm{N}(x\mid m,v)$ the normal density function with mean $m$ and variance $v$ evaluated at $x$, the marginal likelihoods are given by
$p(\hat{\theta}\mid H_{-})$ $=\mathrm{N}(\hat{\theta}\mid\mu,\sigma^{2}+\tau^{2})\times\frac{\Phi(-\mu_{*}/\tau_{*})}{\Phi(-\mu/\tau)}$ (7a)
$p(\hat{\theta}\mid H_{0})$ $=\mathrm{N}(\hat{\theta}\mid 0,\sigma^{2})$ (7b)
$p(\hat{\theta}\mid H_{+})$ $=\mathrm{N}(\hat{\theta}\mid\mu,\sigma^{2}+\tau^{2})\times\frac{\Phi(\mu_{*}/\tau_{*})}{\Phi(\mu/\tau)}$ (7c)
with posterior mean and variance
$\mu_{*}=\frac{\hat{\theta}/\sigma^{2}+\mu/\tau^{2}}{1/\sigma^{2}+1/\tau^{2}}$ and $\tau_{*}^{2}=\frac{1}{1/\sigma^{2}+1/\tau^{2}}.$
Taking ratios of marginal likelihoods produces the Bayes factors
@@
and the Bayes factors for other hypothesis comparisons can be obtained by transitivity and reciprocity, for example, $\mathrm{BF}_{-0}=\mathrm{BF}_{+0}\,/\,\mathrm{BF}_{+-}$. Posterior probabilities can now be obtained by plugging the Bayes factors and prior odds into (1), leading to
[…]
These Bayes factors and posterior probabilities can be monitored as data accumulate to see how the evidence for the hypotheses change. Moreover, they can be used to define the randomization probabilities via (3) leading to
$\pi$ $=\left(1+\frac{\Phi(-\mu_{*}/\tau_{*})}{\Phi(\mu_{*}/\tau_{*})}+\frac{\Pr(H_{0})}{1-\Pr(H_{0})}\,\times\,\exp\left[-\frac{1}{2}\left\{\frac{\hat{\theta}^{2}}{\sigma^{2}}-\frac{(\hat{\theta}-\mu)^{2}}{\sigma^{2}+\tau^{2}}\right\}\right]\,\times\,\frac{\sqrt{1+\tau^{2}/\sigma^{2}}}{\Phi(\mu_{*}/\tau_{*})}\right)^{-1}$
$\phantom{=}+\left(1+\frac{1-\Pr(H_{0})}{\Pr(H_{0})}\,\times\,\exp\left[-\frac{1}{2}\left\{\frac{(\hat{\theta}-\mu)^{2}}{\sigma^{2}+\tau^{2}}-\frac{\hat{\theta}^{2}}{\sigma^{2}}\right\}\right]\,\bigg/\,\sqrt{1+\frac{\tau^{2}}{\sigma^{2}}}\right)^{-1}\,\bigg/\,2.$ (8)
As expected, the randomization probability (3.1) approaches equal randomization as the prior probability of $H_{0}$ increases to one (i.e., $\pi\to 50\%$ as $\Pr(H_{0})\nearrow 1$), while it approaches the ordinary Bayesian posterior tail probability of $\theta>0$ based on a normal prior $\theta\sim\mathrm{N}(\mu,\tau^{2})$ as the prior probability of $H_{0}$ decreases to zero (i.e., $\pi\to\Phi(\mu_{*}/\tau_{*})$ as $\Pr(H_{0})\searrow 0$).
@@
Figure 3: Evolution of Bayesian RAR probabilities for one treatment and control group. In each step one observation from the control and one from the treatment group are simulated from a normal distribution with a true standard deviation of 1 and assuming a true mean difference $\theta=0.25$. Randomization probabilities are then computed assuming a normal spike-and-slab prior centered at zero with standard deviation $\tau=1$, and different prior probabilities $\Pr(H_{0})$.
@@
Figure 3 shows sequences of randomization probabilities computed from simulated normal data. To enable comparison of randomization probabilities across different prior probabilities $\Pr(H_{0})$, the data were not simulated under RAR but by simulating an additional observation from the treatment and control groups at each step. We can see that the probability to randomize to the treatment group based on $\Pr(H_{0})=1$ remains static at $50\%$ (yellow line). In contrast, the randomization probabilities based on $\Pr(H_{0})<1$ tend towards $100\%$ as more data accumulate, as expected, given that the data were simulated under a beneficial treatment effect. Moreover, a clear ordering is visible: Probabilities under $\Pr(H_{0})=0$, which corresponds to Thompson sampling, are the most extreme and may even go strongly in the “wrong” direction. For example, the probability to randomize to control is higher than 75% at some point during the earlier stages of the study. In contrast, probabilities based on $0<\Pr(H_{0})<1$ show the same qualitative behavior but are less extreme. Setting a higher prior probability $\Pr(H_{0})$ thus reduces the variability of randomization probabilities but also makes convergence to a probability of 100% for the more effective treatment slower.
@@ 3.2 More than two groups
Suppose now there are $K>1$ treatment groups and consequently $K$ effect estimates, each estimate quantifying the effect of the corresponding treatment relative to the control. A natural generalization is to stack the estimates into a vector $\boldsymbol{\hat{\theta}}=(\hat{\theta}_{1},\dots,\hat{\theta}_{K})^{\top}$ and assume an approximate $K$-variate normal distribution $\boldsymbol{\hat{\theta}}\mid\boldsymbol{\theta}\sim\mathrm{N}_{K}(\boldsymbol{\theta},\boldsymbol{\boldsymbol{\Sigma}})$, where $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{K})^{\top}$ is the vector of true effects and $\boldsymbol{\boldsymbol{\Sigma}}$ is the covariance matrix of $\boldsymbol{\hat{\theta}}$. For example, $\boldsymbol{\hat{\theta}}$ could be a vector of estimated regression coefficients and $\boldsymbol{\boldsymbol{\Sigma}}$ the corresponding covariance matrix.
The hypotheses from Section 2.2 then translate into
$H_{-}\colon\theta_{i}<0,~i=1,\dots,K$
$H_{0}\colon\theta_{i}=0,~i=1,\dots,K$
$H_{+1}\colon\theta_{1}>0~\text{and}~\theta_{1}>\theta_{i},~i=2,\dots,K$
$\vdots$
$H_{+K}\colon\theta_{K}>0~\text{and}~\theta_{K}>\theta_{i},~i=1,\dots,K-1$
@@
All hypotheses apart from $H_{0}$ are composite and require the specification of a prior distribution. In analogy to the one treatment case, we specify a $K$-variate normal prior $\boldsymbol{\theta}\sim\mathrm{N}_{K}(\boldsymbol{\mu},\boldsymbol{\mathcal{T}})$ and truncate its support to the region of corresponding hypothesis (e.g., for hypothesis $H_{+i}$, the space in $\mathbb{R}^{K}$ where the $i$th component is positive and larger than and all other components). Similarly, the prior hypothesis probabilities $\Pr(H_{+i})$ for $i=1,\dots,K$ may again be specified so that the $\mathrm{N}_{K}(\boldsymbol{\mu},\boldsymbol{\mathcal{T}})$ distribution is recovered when the prior is averaged over the hypotheses.
@@
Figure 4: Illustration of a spike-and-slab prior for a two-dimensional effect $\boldsymbol{\theta}=(\theta_{1},\theta_{2})^{\top}$. A point mass prior at $(0,0)^{\top}$ is assumed under $H_{0}$. A normal prior $\theta\sim\mathrm{N}((0,0)^{\top},\boldsymbol{\mathcal{T}})$ with $\boldsymbol{\mathcal{T}}_{ij}=0.5$ for $i\neq j$ and $\boldsymbol{\mathcal{T}}_{ij}=1$ for $i=j$, with support truncated to the space of the corresponding hypothesis is assumed under $H_{-}$, $H_{+1}$, and $H_{+2}$. The correlation ensures that all treatments receive equal prior probability $\Pr(H_{-})=\Pr(H_{+1})=\Pr(H_{+2})=\{1-\Pr(H_{0})\}/3$.
@@
Figure 4 illustrates such a spike-and-slab prior for the two treatment groups scenario ($K=2$). Specifying a prior covariance matrix $\boldsymbol{\mathcal{T}}$ with uniform correlation of 0.5 ensures that the prior probabilities of all hypotheses but $H_{0}$ are equal, which seems a sensible default. This can also be motivated by the fact that for normal outcomes with a shared control group and equal allocation, mean difference effect estimates are correlated by 0.5 due to the common control group for all treatments.
As in the two-group case (Section 3.1), the normal-normal conjugate framework allows us to derive marginal likelihoods in closed-form. The marginal likelihood under $H_{0}$ is
while the marginal likelihood under $H_{-}$ is
@@
with the first row encoding the constraint of $\theta_{2}$ being positive, and the second and third rows encoding the constraints of $\theta_{2}$ being larger than $\theta_{1}$ and $\theta_{3}$, respectively. As expected, for $K=1$, these marginal likelihoods reduce to (7). Moreover, since they are all available in closed-form, Bayes factors, posterior probabilities, and randomization probabilities are also available in closed-form. Crucially, only the exact evaluation of multivariate normal densities and cumulative distribution functions is required, and both are efficiently implemented in statistical software (e.g., in the mvtnorm R package, Genz and Bretz,, 2009). This thus leads to an efficient Bayesian RAR method that allows experimenters to compute randomization probabilities in complex settings, for instance, multiple regression, where a full Bayesian analysis may involve various complexities, such as priors for nuisance parameters and Markov chain Monte Carlo methods for the computation of posteriors.
Figure 5: Evolution of Bayesian RAR probabilities for 3 treatments and a control group. In each step, one observation from each group is simulated from a normal distribution with a standard deviation of 1 and corresponding mean differences as indicated in the panel titles (treatment 1 is the most effective treatment). Randomization probabilities are computed using a normal spike-and-slab prior centered at the origin and with covariance matrix $\boldsymbol{\mathcal{T}}$ with $\boldsymbol{\mathcal{T}}_{ij}=0.5$ for $i\neq j$ and $\boldsymbol{\mathcal{T}}_{ij}=1$ for $i=j$, and for different prior probabilities $\Pr(H_{0})$.
Figure 5 shows sequences of randomization probabilities computed from simulated normal data with $K=3$ treatment groups. We see again that assigning a prior probability $\Pr(H_{0})=1$ leads to static equal randomization at $\pi_{i}=1/(K+1)=25\%$, while assigning $\Pr(H_{0})=0$ (Thompson sampling) leads to the most variable randomization probabilities. Since data are simulated assuming that treatment 1 is the most effective, randomization probabilities based on $\Pr(H_{0})<1$ converge towards $\pi_{1}=100\%$ and toward $0\%$ for the remaining treatments. While convergence is the fastest for $\Pr(H_{0})=0$, this prior probability also accidentally produces rather high randomization probabilities for the control group at the start of the study, which is less pronounced for positive prior probabilities $\Pr(H_{0})>0$.
4 Point null Bayesian RAR for binary outcomes

### v2

@@ 3 Null hypothesis Bayesian RAR under normality
Suppose that the data are summarized by $y=\{\hat{\theta},\sigma\}$, where $\hat{\theta}$ is an estimate of the true treatment effect $\theta$, and $\sigma$ is the standard error of the estimate. This does not mean that the raw data, from which the estimate is computed, need to be normally distributed. For example, $\hat{\theta}$ could be an estimated (standardized) mean difference, log odds/risk/rate/hazard ratio, risk difference, or regression coefficient. The standard error is often of the form $\sigma=\lambda/\sqrt{n}$, where $n$ is the effective sample size and $\lambda$ the standard deviation of one effective observation. Assume that the estimate is (at least approximately) normally distributed around $\theta$ with its variance $\sigma^{2}$ assumed to be known, i.e., $\hat{\theta}\,|\,\theta\sim\mathrm{N}(\theta,\sigma^{2})$.
@@
If the effect $\theta$ is oriented such that a positive effect indicates treatment benefit, the hypotheses from Section 2 translate into $H_{-}\colon\theta<0$ vs. $H_{0}\colon\theta=0$ vs. $H_{+}\colon\theta>0$. The null hypothesis $H_{0}$ is a simple hypothesis with no free parameters, or equivalently, a point (Dirac) prior at zero. The $H_{-}$ and $H_{+}$ hypotheses are composite hypotheses and require prior distributions for $\theta$. A natural choice is a $\theta\sim\mathrm{N}(\mu,\tau^{2})$ distribution whose support is truncated to the negative and positive side, respectively. In the absence of prior knowledge, it is sensible to center the prior on zero ($\mu=0$) to represent clinical equipoise (Freedman,, 1987), giving equal probability to harmful and beneficial effects. Since $H_{0}$ is a point hypothesis, the prior of $\theta$ under $H_{-}$ and $H_{+}$ must be proper ($\tau<\infty$), as otherwise the posterior probability of $H_{0}$ is 1 for any data $y$. The value of the prior standard deviation $\tau$ must therefore be chosen so that the prior is not too vague (Kass and Raftery,, 1995).
The prior probability of the null hypothesis $\Pr(H_{0})$ represents the a priori plausibility of an absent effect, and also acts as a tuning parameter that controls the variability of RAR. An intuitive default is $\Pr(H_{0})=0.5$, representing equipoise between absent and present effects (Johnson,, 2013). For a given $\Pr(H_{0})$, it is then natural to set the prior probabilities of $H_{+}$ and $H_{-}$ to $\Pr(H_{+})=\{1-\Pr(H_{0})\}\times\Phi(\mu/\tau)$ and $\Pr(H_{-})=\{1-\Pr(H_{0})\}\times\Phi(-\mu/\tau)$, with $\Phi(\cdot)$ the standard normal CDF, so that to the prior distribution averaged over $H_{+}$ and $H_{-}$ is again the $\mathrm{N}(\mu,\tau^{2})$ distribution. For example, specifying $\Pr(H_{0})=0.5$ and a zero-centered prior, we obtain $\Pr(H_{+})=\Pr(H_{-})=0.5\times 0.5=0.25.$ Averaging the prior over all hypotheses leads then to a spike-and-slab prior, as illustrated in Figure 2.
@@
With the priors specified, we can compute the marginal likelihood of the observed effect estimate under each hypothesis and obtain posterior probabilities (1). In the normal likelihood and prior framework, all are available in closed-form. Denoting by $\mathrm{N}(x\,|\,m,v)$ the normal density function with mean $m$ and variance $v$ evaluated at $x$, the marginal likelihoods are
$p(\hat{\theta}\,|\,H_{-})$ $=\mathrm{N}(\hat{\theta}\,|\,\mu,\sigma^{2}+\tau^{2})\times\Phi(-\mu_{*}/\tau_{*})\,/\,\Phi(-\mu/\tau)$
$p(\hat{\theta}\,|\,H_{0})$ $=\mathrm{N}(\hat{\theta}\,|\,0,\sigma^{2})$
$p(\hat{\theta}\,|\,H_{+})$ $=\mathrm{N}(\hat{\theta}\,|\,\mu,\sigma^{2}+\tau^{2})\times\Phi(\mu_{*}/\tau_{*})\,/\,\Phi(\mu/\tau)$
with posterior mean and variance $\mu_{*}=(\hat{\theta}/\sigma^{2}+\mu/\tau^{2})\tau^{2}_{*}$ and $\tau_{*}^{2}=1/(1/\sigma^{2}+1/\tau^{2})$. Taking ratios of marginal likelihoods leads to the Bayes factors
@@
and the Bayes factors for other hypothesis comparisons can be obtained by transitivity and reciprocity (e.g., $\mathrm{BF}_{-0}=\mathrm{BF}_{+0}\,/\,\mathrm{BF}_{+-}$). Posterior probabilities can now be obtained by plugging the Bayes factors and prior odds into (1). Bayes factors and posterior probabilities can be monitored as data accumulate to see how the evidence for the hypotheses changes. Moreover, they can be plugged into (3) to obtain the randomization probability
$\pi$ $=\frac{\overbrace{\Phi\left(\mu_{*}/\tau_{*}\right)}^{\text{Thompson Sampling}}}{1+\frac{\Pr(H_{0})}{1-\Pr(H_{0})}\times\mathrm{BF}_{01}(\hat{\theta})}+\frac{\overbrace{1/2}^{\text{Equal Randomization}}}{1+\frac{1-\Pr(H_{0})}{\Pr(H_{0})}\,\big/\,\mathrm{BF}_{01}(\hat{\theta})}$ (5)
where $\mathrm{BF}_{01}(\hat{\theta})=\exp[-\{(\hat{\theta}/\sigma)^{2}-(\hat{\theta}-\mu)^{2}/(\sigma^{2}+\tau^{2})\}/2]\sqrt{1+\tau^{2}/\sigma^{2}}$ is the Bayes factor contrasting $H_{0}$ to the alternative $H_{1}=H_{-}\cup H_{+}$ with prior $\theta\,|\,H_{1}\sim\mathrm{N}(\mu,\tau^{2})$. As expected, the randomization probability (5) shrinks towards equal randomization as the prior probability of $H_{0}$ increases (i.e., $\pi\to 50\%$ as $\Pr(H_{0})\nearrow 1$), whereas it approaches the ordinary Bayesian posterior tail probability of $\theta>0$ based on a $\theta\sim\mathrm{N}(\mu,\tau^{2})$ prior as the prior probability of $H_{0}$ decreases (i.e., $\pi\to 1-\Phi\{(0-\mu_{*})/\tau_{*}\}=\Phi(\mu_{*}/\tau_{*})$ as $\Pr(H_{0})\searrow 0$). Similarly, the randomization probability shrinks towards 50% as the data provide more evidence for $H_{0}$ ($\pi\to 50\%$ as $\mathrm{BF}_{01}\nearrow\infty$), whereas it approaches the ordinary posterior tail probability as the evidence for $H_{1}$ increases ($\pi\to\Phi(\mu_{*}/\tau_{*})$ as $\mathrm{BF}_{01}\searrow 0$). In contrast to ad hoc modifications of Thompson sampling, shrinkage of the posterior tail probability thus depends directly on the accumulated data via the Bayes factor.
@@
Figure 3: Evolution of Bayesian RAR probabilities for three simulated data sequences (a). In each step, one patient is allocated to the treatment or control group using Thompson sampling assuming a normal prior centered at zero with standard deviation $\tau=1$. Patient outcomes are then simulated from a normal distribution with a true standard deviation of 1 and assuming a true mean difference $\theta$ between treatment and control group as indicated in the plot panels. Randomization probabilities under the spike-and-slab prior ($\Pr(H_{0})>0$) are computed from the same data sequence to enable comparability. Plot (b) shows the distribution of randomization probabilities based on 1’000 simulations under Bayesian RAR for two different prior standard deviations and $\Pr(H_{0})$.
@@
Figure 3a shows sequences of randomization probabilities computed from normal data simulated under different true mean differences between treatment and control group $\theta$ as indicated in the panels. For each $\theta$, one sequence was simulated under RAR using Thompson sampling ($\Pr(H_{0})=0$). RAR probabilities for other values of $\Pr(H_{0})$ were computed from the same sequences to base comparison on the same data. Across all $\theta$, the probability to randomize to the treatment group based on $\Pr(H_{0})=1$ remains static at $50\%$ (yellow). In contrast, for a beneficial treatment effect ($\theta>0$; left plot), the randomization probabilities based on $\Pr(H_{0})<1$ tend towards $100\%$ as more data accumulate, whereas for a harmful treatment effect ($\theta>0$; right plot), they tend towards 0%. Moreover, a clear ordering is visible: Probabilities under $\Pr(H_{0})=0$ (Thompson sampling) are the most extreme both in the “correct” and “wrong” direction. Probabilities based on $0<\Pr(H_{0})<1$ show the same qualitative behavior but are less extreme. Setting a higher $\Pr(H_{0})$ thus reduces the variability of randomization probabilities but also slows down convergence to 100% or 0%. Finally, when there is no treatment effect ($\theta=0$; middle plot), the probabilities based on $0<\Pr(H_{0})<1$ stay relatively close to 50%, whereas the Thompson sampling probabilities tend to wander around.
Figure 3b shows the distribution of randomization probabilities obtained from 1’000 simulations of the scenarios from Figure 3a and for two prior standard deviations $\tau$. When the true means differ from zero ($\theta=\pm 0.25$; first and third row), we see that the randomization probabilities for both $\Pr(H_{0})=0$ (Thompson sampling) and $\Pr(H_{0})=0.5$ correctly converge to 100% and 0%, respectively, although the convergence of $\Pr(H_{0})=0.5$ is slower. In contrast, when there is no difference ($\theta=0$; second row), only the randomization probabilities based on $\Pr(H_{0})=0.5$ converge to 50%, whereas the Thompson sampling randomization probabilities remain roughly uniformly distributed for all sample sizes. The qualitative behavior of the randomization probabilities is the same for both values of the prior standard deviation $\tau$. However, convergence to 100% and 0% is slightly faster for smaller $\tau$ (blue), whereas convergence to 50% is faster for larger $\tau$ (orange).
@@ 3.1 More than two groups
Suppose now there are $K>1$ treatment groups and consequently $K$ effect estimates, each estimate quantifying the effect of the corresponding treatment relative to control. A natural generalization is to stack them into a vector $\boldsymbol{\hat{\theta}}=(\hat{\theta}_{1},\dots,\hat{\theta}_{K})^{\top}$ and assume a $K$-variate normal distribution $\boldsymbol{\hat{\theta}}\,|\,\boldsymbol{\theta}\sim\mathrm{N}_{K}(\boldsymbol{\theta},\boldsymbol{\boldsymbol{\Sigma}})$, where $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{K})^{\top}$ is the vector of true effects and $\boldsymbol{\boldsymbol{\Sigma}}$ is the (assumed to be known) covariance matrix of $\boldsymbol{\hat{\theta}}$. For example, $\boldsymbol{\hat{\theta}}$ could be a vector of estimated regression coefficients and $\boldsymbol{\boldsymbol{\Sigma}}$ its estimated covariance matrix.
@@
The hypotheses from Section 2.1 then translate into $H_{-}\colon\max\{\theta_{1},\dots,\theta_{K}\}<0$ vs. $H_{0}\colon\theta_{1}=\dots=\theta_{K}=0$ vs. $H_{+i}\colon\theta_{i}=\max\{\theta_{1},\dots,\theta_{K}\}>0$ for $i=1,\dots,K$. All hypotheses apart from $H_{0}$ are composite and require the specification of a prior distribution. In analogy to the $K=1$ case, we specify a $K$-variate normal prior $\boldsymbol{\theta}\sim\mathrm{N}_{K}(\boldsymbol{\mu},\boldsymbol{\mathcal{T}})$ and truncate its support to the region of the corresponding hypothesis. Similarly, a sensible default is to center the prior on $\boldsymbol{0}$ to encode clinical equipoise, and specify prior probabilities $\Pr(H_{+i})$ for $i=1,\dots,K$ so that the $\mathrm{N}_{K}(\boldsymbol{\mu},\boldsymbol{\mathcal{T}})$ distribution is recovered when the prior is averaged over the hypotheses.
@@
Figure 4: Illustration of a spike-and-slab prior for a two-dimensional effect $\boldsymbol{\theta}=(\theta_{1},\theta_{2})^{\top}$. A point mass prior at $(0,0)^{\top}$ is assumed under $H_{0}$. A normal prior $\theta\sim\mathrm{N}((0,0)^{\top},\boldsymbol{\mathcal{T}})$ with $\boldsymbol{\mathcal{T}}_{ij}=0.5$ for $i\neq j$ and $\boldsymbol{\mathcal{T}}_{ij}=1$ for $i=j$, with support truncated to the space of the corresponding hypothesis is assumed under $H_{-}$, $H_{+1}$, and $H_{+2}$. The correlation of 0.5 ensures that all treatments receive equal prior probability $\Pr(H_{-})=\Pr(H_{+1})=\Pr(H_{+2})=\{1-\Pr(H_{0})\}/3$.
@@
Figure 4 illustrates a spike-and-slab prior for two treatment groups ($K=2$). The prior covariance matrix $\boldsymbol{\mathcal{T}}$ is set to a uniform correlation of 0.5 which ensures that the prior probabilities of all hypotheses but $H_{0}$ are equal. Note that a zero-centered prior with uniform correlation of 0.5 ensures equal treatment prior probabilities for any number of treatment groups $K$, and thus is a sensible default in the absence of prior knowledege.
@@
Also for $K>1$, marginal likelihoods, Bayes factors, posterior probabilities, and randomization probabilities are available in closed-form (shown in the supplement). Crucially, only the evaluation of multivariate normal densities and CDFs is required. This thus leads to an efficient Bayesian RAR method that allows experimenters to compute randomization probabilities in complex settings, for instance, multiple regression, where a full Bayesian analysis may involve additional complexities, such as priors for nuisance parameters and Markov chain Monte Carlo methods for the computation of posteriors.
4 Null hypothesis Bayesian RAR for binary outcomes